OpenStudy (anonymous):

Can someone explain how the application of this formula: \[x2+px+q=0\] \[x=-p/2 +- \sqrt{(p/2)^2 - q}\] To this problem: \[7x^2 - 5x -2\] Becomes this in factor form: \[7(x-1)(x+2/7)\]

OpenStudy (anonymous):

\[7x^2-5x-2=0\]divide both sides by 7\[x^2-\frac{5}{7}x-\frac{2}{7}=0\]now u can apply that formula

OpenStudy (anonymous):

I tried that but I keep getting it wrong... :/

OpenStudy (anonymous):

what u got ??

OpenStudy (anonymous):

\[x^2-5x/7-2/7\] \[(5x/7)/2 +- \sqrt{((5x/7)/2)^2 + 2/7}\] \[x1: (5x/7)/2 - (5x/7)/2 + \sqrt{2/7} = \sqrt{2/7}\] I should get 2/7 though, not sqrt{2/7}

OpenStudy (anonymous):

dude what did u do ???

OpenStudy (anonymous):

p=-5/7 q=-2/7

OpenStudy (anonymous):

Ehm... wish I knew :)

OpenStudy (anonymous):

q = -(-2/7), isn't it? So +2/7

OpenStudy (anonymous):

\[x=-p/2 +- \sqrt{(p/2)^2 - q}=5/14 +- \sqrt{(5/14)^2 +2/7}\]

OpenStudy (anonymous):

firstly tell me why u dont use\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]??

OpenStudy (anonymous):

because this chapter is about learning the x2+px+q=0 formula

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

although, the quadratic formula seems far better suited to this problem

OpenStudy (anonymous):

u cant seperate radical like that :|

OpenStudy (anonymous):

\[5/14 \pm \sqrt{(5/14)^2 +2/7}=5/14 \pm \sqrt{81/196}=5/14 \pm 9/14\]

OpenStudy (anonymous):

wow... thanks, I can see I'm starting to get tired. This shouldn't be so hard. :/

OpenStudy (anonymous):

yeah :) no problem